OpenStudy (anonymous):

Solve the following equation. 8 + x = 13 + 2x x=

OpenStudy (anonymous):

combine like terms 8+x = 13+2x--- subtract 8 to both sides x = 5+2x --- subtract 2x to both sides -x = 5 --- divide both sides by -1 x = -5